OpenStudy (anonymous):

Solve: 2(x + 1) + 4 = 8 1 6.5 4 3

OpenStudy (anonymous):

Just do: 2(x+1) + 4 = 8 2(x+1) = 8 - 4 2(x+1) = 4 divide by 2 x+1=2 x=1 :)

OpenStudy (whpalmer4):

You can think it out even without algebra. If <something> + 4 is 8, <something> must equal 4, right? If 2<something else> = 4, <something else> must equal 2, right? If <yet another something> + 1 is 2, <yet another something> must be 1, right?